OpenStudy (anonymous):

help please, i cant think straight today. What value of x makes the equation true? –(9 + (–17)) = –9 + x x =

OpenStudy (anonymous):

add 9 to both sides and solve \[-(9+(-17))=-9+x+9\]

OpenStudy (anonymous):

oops sorry forgot to add 9 to the left side

OpenStudy (anonymous):

lol its ok

OpenStudy (anonymous):

now you have x on right side. solve for the left side -(9+(-17))+9 release the brackets: -9 - (-17) + 9=-(-17)=17

OpenStudy (anonymous):

so x equals 17

OpenStudy (anonymous):

oh thank you! :D

OpenStudy (anonymous):

np :) :)

OpenStudy (anonymous):

the reason i cant think straight today is because i got sick :/

OpenStudy (anonymous):

first time i got sick in 4 - 5 years

OpenStudy (anonymous):

well. get better. hope it won't be long. :)

OpenStudy (anonymous):

thank you :) my temperature went from 101.4 to 100.9 so im slowly getting better :)

OpenStudy (anonymous):

:) Get better. :)

OpenStudy (anonymous):

ty :)